# Differentially Private SGD Without Clipping Bias: An Error-Feedback Approach

### v1

@@ Appendix A Proof of results in Section 3
By substituting the above lower bound of $\alpha^{t}_{\mathbf{e}}$ back into the upper bounds of $\eta$, $\beta^{t},\gamma$ in (14), we will get explicit bounds for these constants and $\kappa,\Delta^{t}$. First, it is clear that we can set $\gamma=\frac{1}{3}\leq\frac{1}{2+\alpha^{t}_{\mathbf{e}}}$ because $\alpha^{t}_{\mathbf{e}}\leq 1.$
Next, we derive the range of $\frac{\alpha^{t-1}_{\mathbf{e}}\left(\sqrt{1+16\gamma\frac{1+\alpha^{t-1}_{\mathbf{e}}}{\alpha^{t-1}_{\mathbf{e}}}}-1\right)}{8\gamma(1+\alpha^{t-1}_{\mathbf{e}})L}$ in the upper bound of $\eta$ that $\eta\leq\min\left\{\frac{1}{\mu},\frac{\alpha^{t-1}_{\mathbf{e}}\left(\sqrt{1+16\gamma\frac{1+\alpha^{t-1}_{\mathbf{e}}}{\alpha^{t-1}_{\mathbf{e}}}}-1\right)}{8\gamma(1+\alpha^{t-1}_{\mathbf{e}})L}\right\}.$ We analyze it in two cases: 1) $\alpha^{t-1}_{\mathbf{e}}\leq\frac{1}{2}$ and 2) $\alpha^{t-1}_{\mathbf{e}}>\frac{1}{2}$. In case 1) where $\alpha^{t-1}_{\mathbf{e}}\geq\frac{1}{2}$,
$\frac{\alpha^{t-1}_{\mathbf{e}}\left(\sqrt{1+16\gamma\frac{1+\alpha^{t-1}_{\mathbf{e}}}{\alpha^{t-1}_{\mathbf{e}}}}-1\right)}{8\gamma(1+\alpha^{t-1}_{\mathbf{e}})L}\geq\frac{3}{16L},$
and in case 2) where $\alpha^{t-1}_{\mathbf{e}}<\frac{1}{2},$
$\frac{\alpha^{t-1}_{\mathbf{e}}\left(\sqrt{1+16\gamma\frac{1+\alpha^{t-1}_{\mathbf{e}}}{\alpha^{t-1}_{\mathbf{e}}}}-1\right)}{8\gamma(1+\alpha^{t-1}_{\mathbf{e}})L}>\frac{1.6\sqrt{\alpha^{t-1}_{\mathbf{e}}}}{4L}\geq\frac{2\sqrt{C_{2}}}{5L\sqrt{C_{2}+t(G^{\prime}-C_{2})}}.$
Therefore, we can set $\eta=\min\left\{\frac{1}{2\mu},\frac{3}{32L},\frac{\sqrt{C_{2}}}{5L\sqrt{C_{2}+t(G^{\prime}-C_{2})}}\right\},$ and from (15b), we obtain $\kappa=\min\{\frac{1}{2},\frac{3\mu}{32L},\frac{\mu\sqrt{C_{2}}}{5L\sqrt{C_{2}+t(G^{\prime}-C_{2})}}\}=\eta\mu.$ By substituting the above choices of $\eta,\kappa$ and $\beta^{t}$ into the definition of $\Delta^{t}$ that
$\Delta^{t}$ $=\frac{(1+\gamma\eta L(1+\frac{1}{\beta^{t-1}}))\sigma^{2}}{B}+d\sigma_{1}^{2}+2\gamma\eta L(1+\frac{1}{\beta^{t-1}})C_{1}^{2}$
$\leq\frac{\sigma^{2}}{B}(1+\frac{\sqrt{1+\frac{32}{3}}-1}{16})+d\sigma_{1}^{2}+2\frac{\sqrt{1+\frac{32}{3}}-1}{16}C_{1}^{2}$
$\leq\frac{5\sigma^{2}}{4B}+d\sigma_{1}^{2}+\frac{C_{1}^{2}}{2}.$
Substitute $\eta,\kappa$ and $\Delta^{t}$ into (16), we obtain
@@
$\operatorname{\mathbb{E}}[\mathcal{L}^{T+1}]$ $\leq(1-\mu\eta)^{T}\mathcal{L}^{0}+\frac{\eta L}{2\mu}\left(\frac{5\sigma^{2}}{4B}+d\sigma_{1}^{2}+\frac{C_{1}^{2}}{2}\right).$
The theorem is proved.
@@
Then we start with bounding the update of $\mathbf{e}$ with the following lemma:

### v2

@@ Appendix A Proof of results in Section 3
A.1 Proof of Theorem 3.6
With smoothness Assumption 3.2, we have the following descent property, where the expectation $\operatorname{\mathbb{E}}_{t}[\cdot]$ is taken on the randomness over iteration $t$ conditioned on all past histories from 00 to $t$:
$\operatorname{\mathbb{E}}_{t}$ $[f(\mathbf{x}^{t+1})]\leq f(\mathbf{x}^{t})+\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}_{t}[\mathbf{x}^{t+1}-\mathbf{x}^{t}]\right\rangle+\operatorname{\mathbb{E}}_{t}\left[\frac{L}{2}\left\lVert\mathbf{x}^{t+1}-\mathbf{x}^{t}\right\rVert^{2}\right]$
@@
$\stackrel{{\scriptstyle(a)}}{{\leq}}f(\mathbf{x}^{t})-\eta\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}_{t}[\mathbf{v}^{t}]\right\rangle+\frac{L\eta^{2}}{2}\operatorname{\mathbb{E}}_{t}\left[\left\lVert\mathbf{v}^{t}\right\rVert^{2}\right]+\frac{Ld\eta^{2}}{2}\sigma_{1}^{2}$
$=f(\mathbf{x}^{t})-\eta\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}_{t}[\mathbf{v}^{t}]\right\rangle+\frac{L\eta^{2}}{2}\operatorname{\mathbb{E}}_{t}\left[\left\lVert\frac{1}{B}\sum_{i\in\mathcal{B}^{t}}\mathrm{clip}\left(\mathbf{g}^{t}_{i},C_{1}\right)+\mathrm{clip}\left(\mathbf{e}^{t},C_{2}\right)\right\rVert^{2}\right]+\frac{Ld\eta^{2}}{2}\sigma_{1}^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}f(\mathbf{x}^{t})-\eta\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}_{t}[\mathbf{v}^{t}]\right\rangle+\frac{L\eta^{2}}{2}(C_{1}+C_{2})^{2}+\frac{Ld\eta^{2}}{2}\sigma_{1}^{2},$ (10)
where $(a)$ applies the update rule of $\mathbf{x}^{t}$ and uses the fact that $\mathbf{w}^{t}$ follows zero-mean Gaussian and is independent of $\mathbf{v}^{t}$, and $(b)$ bounds $\left\lVert\mathbf{v}^{t}\right\rVert$ by $C_{1}+C_{2}$ with clipping operation. Next, we bound the inner-product between $\nabla f(\mathbf{x}^{t})$ and $\operatorname{\mathbb{E}}[\mathbf{v}^{t}]$ as:
[…]
$=-\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}\alpha^{t}_{i}\mathbf{g}^{t}_{i}+\operatorname{\mathbb{E}}\alpha_{\mathbf{e}}^{t}\mathbf{e}^{t}\right\rangle$
$=-\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}\alpha^{t}_{i}\mathbf{g}^{t}_{i}\right\rangle-\left\langle\nabla f(\mathbf{x}^{t})-\nabla f(\mathbf{x}^{t-1}),\operatorname{\mathbb{E}}\alpha_{\mathbf{e}}^{t}\mathbf{e}^{t}\right\rangle-\left\langle\nabla f(\mathbf{x}^{t-1}),\operatorname{\mathbb{E}}\alpha_{\mathbf{e}}^{t}\mathbf{e}^{t}\right\rangle$
$\stackrel{{\scriptstyle(a)}}{{\leq}}-\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}\alpha^{t}_{i}\mathbf{g}^{t}_{i}\right\rangle+\frac{1}{2}\left(\frac{1}{\eta L}\operatorname{\mathbb{E}}\left\lVert\nabla f(\mathbf{x}^{t})-\nabla f(\mathbf{x}^{t-1})\right\rVert^{2}+\eta L\left\lVert\alpha_{\mathbf{e}}^{t}\mathbf{e}^{t}\right\rVert^{2}\right)$
$\quad-\left\langle\nabla f(\mathbf{x}^{t-1}),\operatorname{\mathbb{E}}\alpha_{\mathbf{e}}^{t}\left((1-\alpha^{t-1}_{\mathbf{e}})\mathbf{e}^{t-1}+\nabla f(\mathbf{x}^{t-1})-\frac{1}{N}\sum^{N}_{i=1}\alpha^{t-1}_{i}\mathbf{g}^{t-1}_{i}\right)\right\rangle$
$\stackrel{{\scriptstyle(b)}}{{\leq}}-\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}\alpha^{t}_{i}\mathbf{g}^{t}_{i}\right\rangle+\frac{\eta L}{2}\left(\operatorname{\mathbb{E}}\left\lVert\mathbf{v}^{t-1}+\mathbf{w}^{t-1}\right\rVert^{2}+\operatorname{\mathbb{E}}\left\lVert\alpha_{\mathbf{e}}^{t}\mathbf{e}^{t}\right\rVert^{2}\right)$
$\quad-\frac{\alpha_{\mathbf{e}}^{t}(1-\alpha^{t-1}_{\mathbf{e}})}{\alpha^{t-1}_{\mathbf{e}}}\left\langle\nabla f(\mathbf{x}^{t-1}),\operatorname{\mathbb{E}}\alpha_{\mathbf{e}}^{t-1}\mathbf{e}^{t-1}\right\rangle-\alpha_{\mathbf{e}}^{t}\left\lVert\nabla f(\mathbf{x}^{t-1})\right\rVert^{2}+\alpha_{\mathbf{e}}^{t}\left\langle\nabla f(\mathbf{x}^{t-1}),\operatorname{\mathbb{E}}\alpha^{t-1}_{i}\mathbf{g}^{t-1}_{i}\right\rangle$
$\stackrel{{\scriptstyle(c)}}{{\leq}}-\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}\alpha^{t}_{i}\mathbf{g}^{t}_{i}\right\rangle+\frac{\eta L}{2}\left((C_{1}+C_{2})^{2}+d\sigma_{1}^{2}+C_{2}^{2}\right)-\alpha_{\mathbf{e}}^{t}\left\lVert\nabla f(\mathbf{x}^{t-1})\right\rVert^{2}$
$\quad-\frac{\alpha_{\mathbf{e}}^{t}(1-\alpha^{t-1}_{\mathbf{e}})}{\alpha^{t-1}_{\mathbf{e}}}\left\langle\nabla f(\mathbf{x}^{t-1}),\operatorname{\mathbb{E}}\alpha_{\mathbf{e}}^{t-1}\mathbf{e}^{t-1}\right\rangle+\alpha_{\mathbf{e}}^{t}\left\langle\nabla f(\mathbf{x}^{t-1}),\operatorname{\mathbb{E}}\alpha^{t-1}_{i}\mathbf{g}^{t-1}_{i}\right\rangle,$
where $(a)$ applies (6) to the second term and expand $\mathbf{e}^{t}$ with (8); $(b)$ applies the smoothness Assumption 3.2 to the second term; $(c)$uses the fact that $\mathbf{w}^{t-1}$ follows zero-mean Gaussian and is independent of $\mathbf{v}^{t-1}$ and that $\mathbf{v}^{t-1}=\mathrm{clip}\left(\mathbf{e}^{t-1},C_{2}\right)+\frac{1}{B}\sum_{i\in\mathcal{B}^{t-1}}\mathrm{clip}\left(\mathbf{g}^{t-1}_{i},C_{1}\right)$ has magnitude less or equal to $C_{1}+C_{2}$.
By recursively expanding the term $-\left\langle\nabla f(\mathbf{x}^{\tau}),\operatorname{\mathbb{E}}\alpha^{\tau}_{\mathbf{e}}\mathbf{e}^{\tau}\right\rangle$ above as
$-\left\langle\nabla f(\mathbf{x}^{\tau}),\operatorname{\mathbb{E}}\alpha^{\tau}_{\mathbf{e}}\mathbf{e}^{\tau}\right\rangle\leq\frac{\eta L}{2}\left((C_{1}+C_{2})^{2}+d\sigma_{1}^{2}+C_{2}^{2}\right)-\alpha_{\mathbf{e}}^{\tau}\left\lVert\nabla f(\mathbf{x}^{\tau-1})\right\rVert^{2}$
$\quad-\frac{\alpha_{\mathbf{e}}^{t}(1-\alpha^{\tau-1}_{\mathbf{e}})}{\alpha^{\tau-1}_{\mathbf{e}}}\left\langle\nabla f(\mathbf{x}^{\tau-1}),\operatorname{\mathbb{E}}\alpha_{\mathbf{e}}^{\tau-1}\mathbf{e}^{\tau-1}\right\rangle+\alpha_{\mathbf{e}}^{\tau}\left\langle\nabla f(\mathbf{x}^{\tau-1}),\operatorname{\mathbb{E}}_{\tau-1}\alpha^{\tau-1}_{i}\mathbf{g}^{\tau-1}_{i}\right\rangle,$
and note that $\mathbf{e}^{0}=0,\alpha_{\mathbf{e}}^{0}=1$, we have:
-- $\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}[\mathbf{v}^{t}]\right\rangle=-\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}\alpha^{t}_{i}\mathbf{g}^{t}_{i}\right\rangle-\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}\alpha_{\mathbf{e}}^{t}\mathbf{e}^{t}\right\rangle$
$\leq-\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}\alpha^{t}_{i}\mathbf{g}^{t}_{i}\right\rangle+\frac{\eta L}{2}\left((C_{1}+C_{2})^{2}+d\sigma_{1}^{2}+C_{2}^{2}\right)-\alpha_{\mathbf{e}}^{t}\left\lVert\nabla f(\mathbf{x}^{t-1})\right\rVert^{2}$
$\quad-\frac{\alpha_{\mathbf{e}}^{t}(1-\alpha^{t-1}_{\mathbf{e}})}{\alpha^{t-1}_{\mathbf{e}}}\left\langle\nabla f(\mathbf{x}^{t-1}),\operatorname{\mathbb{E}}\alpha_{\mathbf{e}}^{t-1}\mathbf{e}^{t-1}\right\rangle+\alpha_{\mathbf{e}}^{t}\left\langle\nabla f(\mathbf{x}^{t-1}),\operatorname{\mathbb{E}}\alpha^{t-1}_{i}\mathbf{g}^{t-1}_{i}\right\rangle$
$\leq-\sum^{t-1}_{\tau=0}\alpha_{\mathbf{e}}^{t}\left(\prod^{t-1}_{\tau_{1}=\tau+1}(1-\alpha^{\tau_{1}}_{\mathbf{e}})\right)\left\lVert\nabla f(\mathbf{x}^{\tau})\right\rVert^{2}-\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}\alpha^{t}_{i}\mathbf{g}^{t}_{i}\right\rangle$
$\quad+\sum^{t}_{\tau=1}\alpha^{t}_{\mathbf{e}}\left(\prod^{t-1}_{\tau_{1}=\tau}(1-\alpha^{\tau_{1}}_{\mathbf{e}})\right)\left(\frac{L\eta}{2}\left(2C_{1}^{2}+3C_{2}^{2}+d\sigma_{1}^{2}\right)+\left\langle\nabla f(\mathbf{x}^{\tau}),\operatorname{\mathbb{E}}\alpha_{i}^{\tau}\mathbf{g}^{\tau}_{i}\right\rangle\right).$ (11)
Substitute (A.1) back to (A.1) and sum over iterations, we have:
$\operatorname{\mathbb{E}}$ $[f(\mathbf{x}^{T})]\leq f(\mathbf{x}^{0})-\eta\sum^{T-1}_{t=0}\sum^{t-1}_{\tau=0}\alpha_{\mathbf{e}}^{t}\left(\prod^{t-1}_{\tau_{1}=\tau+1}(1-\alpha^{\tau_{1}}_{\mathbf{e}})\right)\left\lVert\nabla f(\mathbf{x}^{\tau})\right\rVert^{2}$
$\quad+\frac{L\eta^{2}}{2}\sum^{T-1}_{t=0}\left(\sum^{t}_{\tau=1}\alpha^{t}_{\mathbf{e}}\left(\prod^{t-1}_{\tau_{1}=\tau}(1-\alpha^{\tau_{1}}_{\mathbf{e}})\right)\left(2C_{1}^{2}+3C_{2}^{2}+d\sigma_{1}^{2}\right)+2C_{1}^{2}+2C_{2}^{2}+d\sigma_{1}^{2}\right)$
$\quad-\eta\sum^{T-1}_{t=0}\left(\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}\alpha_{i}^{t}\mathbf{g}^{t}_{i}\right\rangle-\sum^{t}_{\tau=1}\alpha^{t}_{\mathbf{e}}\left(\prod^{t-1}_{\tau_{1}=\tau}(1-\alpha^{\tau_{1}}_{\mathbf{e}})\right)\left\langle\nabla f(\mathbf{x}^{\tau}),\operatorname{\mathbb{E}}\alpha_{i}^{\tau}\mathbf{g}^{\tau}_{i}\right\rangle\right)$
=$\stackrel{{\scriptstyle(a)}}{{=}}f(\mathbf{x}^{0})-\eta\sum^{T-1}_{t=0}\sum^{T-1}_{\tau=t+1}\alpha_{\mathbf{e}}^{\tau}\left(\prod^{\tau-1}_{\tau_{1}=t+1}(1-\alpha^{\tau_{1}}_{\mathbf{e}})\right)\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}$
$\quad+\frac{L\eta^{2}}{2}\sum^{T-1}_{t=0}\left(\sum^{t}_{\tau=1}\alpha^{t}_{\mathbf{e}}\left(\prod^{t-1}_{\tau_{1}=\tau}(1-\alpha^{\tau_{1}}_{\mathbf{e}})\right)\left(2C_{1}^{2}+3C_{2}^{2}+d\sigma_{1}^{2}\right)+2C_{1}^{2}+2C_{2}^{2}+d\sigma_{1}^{2}\right)$
$\quad-\eta\sum^{T-1}_{t=0}\left(1-\sum^{T-1}_{\tau=t+1}\alpha^{\tau}_{\mathbf{e}}\left(\prod^{\tau-1}_{\tau_{1}=t+1}(1-\alpha^{\tau_{1}}_{\mathbf{e}})\right)\right)\left\langle\nabla f(\mathbf{x}^{t}),\operatorname{\mathbb{E}}\alpha_{i}^{t}\mathbf{g}^{t}_{i}\right\rangle$
$\stackrel{{\scriptstyle(b)}}{{\leq}}f(\mathbf{x}^{0})-\eta\sum^{T-1}_{t=0}A_{t}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}+\frac{L\eta^{2}}{2}B_{T}(2C_{1}^{2}+3C_{2}^{2}+d\sigma_{1}^{2})+\eta\sum^{T-1}_{t=0}C_{t}C_{1}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert,$
where in $(a)$ we rearrange the terms; in $(b)$ we apply Cauchy-Schwarz inequality to the last term and define $A_{t},B_{T},C_{t}$ accordingly. Specifically, we have
$A_{t}$ $:=\sum^{T-1}_{\tau=t+1}\alpha_{\mathbf{e}}^{\tau}\left(\prod^{\tau-1}_{\tau_{1}=t+1}(1-\alpha^{\tau_{1}}_{\mathbf{e}})\right),$ (12)
$B_{T}$ $:=T+\sum^{T-1}_{t=0}\sum^{t}_{\tau=1}\alpha^{t}_{\mathbf{e}}\left(\prod^{t-1}_{\tau_{1}=\tau}(1-\alpha^{\tau_{1}}_{\mathbf{e}})\right)=T+\sum^{T-1}_{t=0}A_{t},$ (13)
$C_{t}$ $:=\left\lvert 1-\sum^{T-1}_{\tau=t+1}\alpha^{\tau}_{\mathbf{e}}\left(\prod^{\tau-1}_{\tau_{1}=t+1}(1-\alpha^{\tau_{1}}_{\mathbf{e}})\right)\right\rvert=\left\lvert 1-A_{t}\right\rvert.$ (14)
Next, we bound $A_{t},B_{T},C_{t}$, respectively. To begin with, we provide the upper and lower bounds on $A_{t}$. First, note that by definition of $\alpha^{t}_{\mathbf{e}}$ in (8), we have $\alpha^{t}_{\mathbf{e}}\in(0,1]$, therefore the product $\prod^{\tau-1}_{\tau_{1}=t+1}(1-\alpha^{\tau_{1}}_{\mathbf{e}})$ is strictly less than $1$. Let $T^{\prime}\in\{0,\dots T-1\}$ denote the iteration where $\alpha^{t}_{\mathbf{e}}<1,\forall t>T^{\prime}$ and $\alpha^{T^{\prime}}_{\mathbf{e}}=1$. Then $A_{t}$ can be expressed as:
$A_{t}$ $=\sum^{T-1}_{\tau=t+1}\alpha_{\mathbf{e}}^{\tau}\left(\prod^{\tau-1}_{\tau_{1}=t+1}(1-\alpha^{\tau_{1}}_{\mathbf{e}})\right)$
$=1-1+\alpha^{t+1}_{\mathbf{e}}+\alpha^{t+2}_{\mathbf{e}}(1-\alpha^{t+1}_{\mathbf{e}})+\dots+\alpha^{T-1}_{\mathbf{e}}\prod^{T-2}_{\tau={t+1}}(1-\alpha^{\tau}_{\mathbf{e}})$
$=1-(1-\alpha^{t+1}_{\mathbf{e}})+\alpha^{t+2}_{\mathbf{e}}(1-\alpha^{t+1}_{\mathbf{e}})+\dots+\alpha^{T-1}_{\mathbf{e}}\prod^{T-2}_{\tau=t+1}(1-\alpha^{\tau}_{\mathbf{e}})$
$=1-\prod^{T-1}_{\tau=t+1}(1-\alpha^{\tau}_{\mathbf{e}})=\begin{cases}1,&t<T^{\prime},\\ 1-\prod^{T-1}_{\tau=t+1}(1-\alpha^{\tau}_{\mathbf{e}}),&t\geq T^{\prime}.\end{cases}$
Therefore, $B_{T}\leq 2T$. $C_{t}=1-A_{t}\in[0,1)$. Next, we show that the following relation holds: Case I: $t<T^{\prime},A_{t}=1$, it is clear that
$A_{t}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}\geq(1-A_{t})C_{1}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert=0.$
Case II: $t\geq T^{\prime}$, then $A_{t}<1$, we want to show the following relation holds:
$\alpha^{t+1}_{\mathbf{e}}\sum^{t}_{\tau=T^{\prime}}\prod^{t}_{\tau_{1}=\tau+1}(1-\alpha_{\mathbf{e}}^{\tau_{1}})\left\lVert\nabla f(\mathbf{x}^{\tau})\right\rVert^{2}>(1-A_{t})C_{1}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert.$
Let us consider the worst case where $t=T^{\prime}$, i.e., when the left-hand-side only has one term $\alpha^{t+1}_{\mathbf{e}}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}$, then, we have
$\alpha^{t+1}_{\mathbf{e}}=\frac{C_{2}}{\left\lVert\mathbf{e}^{t+1}\right\rVert}$ $<1$
$\frac{C_{2}}{\left\lVert\mathbf{e}^{t}-\alpha^{t}_{\mathbf{e}}\mathbf{e}^{t}+\frac{1}{B}\sum_{i\in\mathcal{B}^{t}}(\mathbf{g}^{t}_{i}-\alpha_{i}^{t}\mathbf{g}^{t}_{i})\right\rVert}$ <(8)${{<}}1$
$C_{2}$ <$\stackrel{{\scriptstyle(a)}}{{<}}\left\lVert\mathbf{e}^{t}-1\cdot\mathbf{e}^{t}+\frac{1}{B}\sum_{i\in\mathcal{B}^{t}}(\mathbf{g}^{t}_{i}-\alpha_{i}^{t}\mathbf{g}^{t}_{i})\right\rVert$
$C_{2}$ <$\stackrel{{\scriptstyle(b)}}{{<}}\left\lVert\nabla f(\mathbf{x}^{t})+\frac{1}{B}\sum_{i\in\mathcal{B}^{t}}(\mathbf{g}^{t}_{i}-\nabla f(\mathbf{x}^{t})-\alpha_{i}^{t}\mathbf{g}^{t}_{i})\right\rVert$
$C_{2}$ <$\stackrel{{\scriptstyle(c)}}{{<}}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert+\frac{\sigma}{B}+\frac{1}{B}\sum_{i\in\mathcal{B}^{t}}\left\lVert\alpha_{i}^{t}\mathbf{g}^{t}_{i}\right\rVert$
$C_{2}-\frac{\sigma}{B}-C_{1}$ <$\stackrel{{\scriptstyle(d)}}{{<}}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert,$ (15)
where $(a)$ uses the fact that $\alpha^{t}_{\mathbf{e}}=1$ at $t=T^{\prime}$; $(b)$ we add and subtract $\nabla f(\mathbf{x}^{t})$; $(c)$ applies triangle inequality and Assumption 3.4; and $(d)$ we arrange the terms and use the fact that $\left\lVert\alpha_{i}^{t}\mathbf{g}^{t}_{i}\right\rVert\leq C_{1}$. By setting $C_{2}\geq 3C_{1}+\frac{\sigma}{B}$, we have $\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert\geq 2C_{1}$. Further, from (15), we also have that
$\alpha^{t+1}_{\mathbf{e}}$ $=\frac{C_{2}}{\left\lVert\mathbf{e}^{t}-\alpha^{t}_{\mathbf{e}}\mathbf{e}^{t}+\frac{1}{B}\sum_{i\in\mathcal{B}^{t}}(\mathbf{g}^{t}_{i}-\alpha_{i}^{t}\mathbf{g}^{t}_{i})\right\rVert}$
$\geq\frac{C_{2}}{\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert+\frac{\sigma}{B}+C_{1}},$ (16)
where we apply triangle inequality to the denominator in the second inequality. Therefore, we have for the worst case:
$\alpha^{t+1}_{\mathbf{e}}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}$ (16)${{\geq}}\frac{C_{2}}{\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert+C_{1}+\frac{\sigma}{B}}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert\cdot\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert$ ) ∥
$\stackrel{{\scriptstyle(a)}}{{\geq}}\frac{(3C_{1}+\frac{\sigma}{B})2C_{1}}{3C_{1}+\frac{\sigma}{B}}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert\geq 2C_{1}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert>2(1-A_{t})C_{1}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert,$
where $(a)$ uses the fact that $\frac{C_{2}}{\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert+C_{1}+\frac{\sigma}{B}}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert$ is monotonically increasing w.r.t. $\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert$. When $t>T^{\prime}$, a similar proof can be applied to show that $\alpha^{t+1}_{\mathbf{e}}\sum^{t}_{\tau=T^{\prime}}\prod^{t}_{\tau_{1}=\tau+1}(1-\alpha_{\mathbf{e}}^{\tau_{1}})\left\lVert\nabla f(\mathbf{x}^{\tau})\right\rVert\geq 2C_{1}.$ Putting the above results together, we have:
$\operatorname{\mathbb{E}}[f(\mathbf{x}^{T})]$ $\leq f(\mathbf{x}^{0})-\frac{\eta}{2}\sum^{T-1}_{t=0}A_{t}\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}+\frac{TL\eta^{2}}{2}(2C_{1}^{2}+3C_{2}^{2}+d\sigma_{1}^{2}),$ (17)
$\operatorname{\mathbb{E}}_{t}[\left\lVert\nabla f(\mathbf{x}^{t})\right\rVert^{2}]$ $\leq\frac{2(f(\mathbf{x}^{0})-f^{\star})}{\eta T}+\eta L(2C_{1}^{2}+3C_{2}^{2}+d\sigma_{1}^{2}),$
where $C_{2}\geq 3C_{1}+\frac{\sigma}{B}$, and the expectation is taken over $t\in\{0,\dots,T-1\}$, with probability $A_{t}/\sum^{T-1}_{t=0}A_{t}$. The theorem is proved.
@@
Then, we start with bounding the update of $\mathbf{e}$ with the following lemma: